### v1

# Communication-Efficient and Tensorized Federated Fine-Tuning of Large Language Models

Abstract
Parameter-efficient fine-tuning (PEFT) methods typically assume that Large Language Models (LLMs) are trained on data from a single device or client. However, real-world scenarios often require fine-tuning these models on private data distributed across multiple devices. Federated Learning (FL) offers an appealing solution by preserving user privacy, as sensitive data remains on local devices during training. Nonetheless, integrating PEFT methods into FL introduces two main challenges: communication overhead and data heterogeneity. In this paper, we introduce FedTT and FedTT+, methods for adapting LLMs by integrating tensorized adapters into client-side models’ encoder/decoder blocks. FedTT is versatile and can be applied to both cross-silo FL and large-scale cross-device FL. FedTT+, an extension of FedTT tailored for cross-silo FL, enhances robustness against data heterogeneity by adaptively freezing portions of tensor factors, further reducing the number of trainable parameters. Experiments on BERT and LLaMA models demonstrate that our proposed methods successfully address data heterogeneity challenges and perform on par or even better than existing federated PEFT approaches while achieving up to 10$\times$ reduction in communication cost.
1 Introduction
Large Language Models (LLMs) like GPT-4 [1], LLaMA [2], and BERT [3] have become transformative tools in AI, excelling at tasks like translation and summarization [4], thanks to their advanced architectures such as Transformers [5]. These models can also be fine-tuned for specialized tasks using task-specific data [6], making them highly adaptable for various applications. However, due to the billions of parameters in modern LLMs, fully fine-tuning these models is computationally expensive and risks overfitting. To mitigate this, parameter-efficient fine-tuning (PEFT) methods like Adapters [7], Prompt-Tuning [8], and LoRA [9] have been developed, optimizing only a subset of parameters while freezing the majority, thus reducing computational costs without sacrificing performance [10].
While PEFT methods typically assume LLMs are trained on data from a single machine or client, real-world applications often require fine-tuning on private data distributed across multiple devices, such as sensitive medical records or legal documents [11, 12]. Privacy concerns make central data collection infeasible, highlighting the need for fine-tuning techniques that can adapt LLMs at the edge while preserving privacy. In this context, Federated Learning (FL) emerges as a compelling solution, as it prioritizes user data privacy by ensuring sensitive information remains on individual devices during training. Instead of transmitting data to a central server, clients in FL locally update their model parameters and share only model information, such as parameters or gradients, which are then aggregated by the server into a global model [13].
To address the aforementioned problems, FL has been integrated into PEFT methods  [14, 15, 16]. These methods, however, often result in two major challenges: (i) communication and computation overhead in the FL system, and (ii) significant accuracy degradation, particularly under heterogeneous scenarios.
Communication efficiency is crucial in FL, as edge devices typically have limited storage and computational power. While some methods attempt to reduce the number of training parameters by incorporating sparsity in PEFT approaches, they either perform poorly in FL [17, 18] or are computationally expensive for clients [19], rendering them unsuitable for practical scenarios.
Data heterogeneity happens when training data is not identically and independently distributed across clients (non-i.i.d.). In such scenarios, local models on individual clients can diverge from the global model’s optimal state, leading to slower convergence [20, 21]. This issue is particularly pronounced when training LLMs in federated environments, as data heterogeneity can severely impact model performance [14]. Although various studies have proposed methods to address this challenge in federated learning [22, 23], these approaches often become inefficient in terms of communication and computational demands when applied to federated LLMs. Several works have modified LoRA to improve its efficiency in highly heterogeneous federated settings [24, 25]. However, these methods require communicating a large number of parameters (at least as much as LoRA).
In this work, we present a federated and tensorized framework that reduces communication load compared to other PEFT methods, while maintaining comparable or even superior accuracy, and addressing challenges such as data heterogeneity. We introduce the Federated Tensor Train (FedTT) algorithm, where tensorized adapters serve as trainable parameters embedded within the encoder/decoder blocks of models used by clients. FedTT is applicable to both cross-silo FL, where all clients participate in training, and large-scale cross-device FL, where only a subset of clients is selected in each round. For cross-silo FL, we propose the FedTT+ algorithm, a heterogeneity-robust method that further reduces the number of parameters by adaptively freezing portions of the tensor factors in the models. We summarize our contributions as follows:
  • We study the TT decomposition of adapters to enable communication-efficient federated fine-tuning of LLMs.
  • We propose FedTT, an approach that effectively fine-tunes LLMs using tensorized adapters, achieving up to 10×10$\times$10 × less communication overhead compared to other popular federated PEFT methods.
  • We introduce FedTT+, an enhancement of FedTT that further reduces the number of trainable parameters and improves robustness against data heterogeneity in cross-silo FL. We demonstrate that FedTT+ outperforms state-of-the-art (SOTA) cross-silo FL methods with fewer trainable parameters.
  • We conduct extensive experiments with widely-used LLMs including the BERT and LLaMA-2 model families to validate the effectiveness of the proposed algorithms in reducing communication load while maintaining high accuracy.
2 Related Work
2.1 Parameter Efficient Fine-Tuning (PEFT)
PEFT methods can be generally divided into three categories [26]. The first is Additive PEFT, where a small set of trainable parameters is added to the model, and only these are updated during training. Methods like Serial Adapter [7], Parallel Adapter [27], Prefix-tuning [28], and Prompt-tuning [8] fall into this category, and our approach aligns with this strategy. The second is Selective PEFT, which chooses a subset of existing model parameters for tuning, as seen in techniques like BitFit [29], U-Diff pruning [30], and PaFi [31]. Lastly, Reparameterized PEFT introduces a low-rank parameterization of pre-trained weights for training, with methods such as LoRA [9] and DoRA [32].
2.2 PEFT in Federated Setting
[14, 15] test and compare various PEFT methods such as Adapter, LoRA, Prompt Tuning, and BitFit in FL. Several works have modified LoRA to improve its efficiency in highly heterogeneous federated settings. For instance, SLoRA [24, 25] modify the initialization to address data heterogeneity, whereas HetLoRA [33] and FlexLoRA [34] adaptively adjust LoRA ranks for each client to handle system heterogeneity. In a recent work, FLoRA [35] slack LoRA matrices $A$ and $B$ for all clients and multiply two resulted matrices to prevent the interference resulting from FedAvg algorithm. However, these methods require communicating a large number of parameters (at least as much as LoRA).
Recently, [19] introduced sparse fine-tuning to reduce the communication load in federated LoRA. While this method lowers communication overhead and shows robustness to data heterogeneity in some tasks, it suffers from computational inefficiency due to the extensive matrix computations required during each communication round, both on the server and client sides. Additionally, their algorithm and results are limited to large-scale cross-device FL scenarios. FFA-LoRA [36] and RoLoRA [37] aim to improve accuracy in the presence of heterogeneity while simultaneously reducing the number of trainable parameters. Our algorithms are not only efficient in terms of communication overhead but also achieve comparable or better accuracy to these approaches particularly when data heterogeneity exists.
2.3 Tensor-based Model Compression
Recent advances in tensor compression have highlighted its potential to reduce model size and enhance training efficiency. Early work by novikov2015tensorizing [38] introduced the TT format, which represents weight matrices as sequences of tensor factors, laying the groundwork for subsequent developments. Recent efforts in tensorized fine-tuning of LLMs have shown promise in achieving high performance with significantly fewer parameters compared to traditional fine-tuning methods. However, these approaches still involve training a large number of variables, especially when compared to the more recent methods like LoRA. In response, yang2024loretta [39] developed the Low-Rank Economic Tensor-Train Adaptation (LoRETTA), which innovates by using tensorized-layer based adapters (LoRETTAadp) and reshaping update matrices into smaller tensor factors (LoRETTArep). Despite demonstrating substantial reductions in trainable parameters—up to 100 times less than popular PEFT methods—the performance of these tensorized approaches in FL scenarios, particularly in the presence of data heterogeneity, remains an open question.
3 Preliminaries
Figure 1: Illustration for the tensorized linear layer (a), FedTT algorithm (b), and the optional tensorized classifier applied for classification tasks (c). Note that we keep the model’s head uncompressed for language modeling tasks. The FedTT algorithm workflow includes fine-tuning clients’ tensorized adapters, aggregating tensor factors on the server, and broadcasting the updated weights back to clients.
3.1 Federated Fine-tuning
Federated fine-tuning is a distributed approach for collaboratively fine-tuning a global model across a central server and a network of $N$ clients, denoted by $\mathcal{C}=\{c_{1},\ldots,c_{N}\}$. The objective is to optimize the global trainable parameters $\bm{w}$ by minimizing the following objective function:
$\min_{\bm{w}}f(\tilde{\bm{w}},\bm{w})=\frac{1}{N}\sum_{i=1}^{N}f_{i}(\tilde{\bm{w}},\bm{w}),$
where $\tilde{\bm{w}}$ is the pre-trained model parameters, which are identical and fixed for all clients, and $f_{i}(\cdot)$ represents the local objective, expressed as:
$f_{i}(\tilde{\bm{w}},\bm{w})=\ell_{i}(\tilde{\bm{w}},\bm{w};\mathcal{D}_{i}),$
where $\ell_{i}(\cdot)$ is the local loss function, and $\mathcal{D}_{i}$ represents the local data distribution for client $c_{i}$.
A well-known FL method is the FedAvg algorithm [13]. The server selects a subset of clients $\mathcal{S}\subseteq\mathcal{C}$ in each training round. Each chosen client $c_{s}\in\mathcal{S}$ initializes its local model with the global model parameters from the previous round $\bm{w}^{(t)}$, and then performs local training using stochastic gradient descent on its dataset for $K$ local updates:
$\bm{w}_{s}^{(t)+k+1}\longleftarrow\bm{w}_{s}^{(t)+k}-\eta\nabla f_{s}(\tilde{\bm{w}},\bm{w}_{s}^{(t)+k}),$
where $\eta$ is the learning rate and $\bm{w}_{s}^{(t)+k}$ refers to the local model parameters for client $c_{s}$ during communication round $t$ and local update $k$. After local training is completed, clients send their updated model parameters to the server, which then aggregates these updates to form the new global model:
$\boldsymbol{w}^{(t+1)}=\frac{1}{|\mathcal{S}|}\sum_{s=1}^{|\mathcal{S}|}\boldsymbol{w}_{s}^{(t)+K}.$
Despite its advantages, FL faces two major challenges: (i) the large size of local models, which results in significant communication overhead, and (ii) data heterogeneity, which can cause local models to diverge from each other (client drift).
3.2 Tensor-Train Decomposition
In this subsection, we provide some introduction to the tensor and tensor train (TT) decomposition [40]. Tensors are natural multidimensional generalizations of matrices. The tensor $\mathcal{W}\in\mathbb{R}^{k_{1}\times\cdots\times k_{J}}$ is indexed as $\mathcal{W}=(w_{i_{1}\cdots i_{J}})_{1\leq i_{j}\leq k_{j}}$ said to have order of $J$ and dimension $k_{1},\cdots,k_{J}$. Given two tensors $\mathcal{W}\in\mathbb{R}^{k_{1}\times\cdots\times k_{J}}$ and $\mathcal{V}\in\mathbb{R}^{l_{1}\times\cdots\times l_{M}}$ with $k_{s}=l_{t}$, the multiplication between two tensors $\mathcal{C}=\mathcal{W}\times_{s,t}\mathcal{V}$ can be performed as:
$\mathcal{C}_{(i_{p})_{p\neq s},(j_{p})_{p\neq t}}=\sum_{i_{s}=j_{t}=1}^{k_{s}}w_{i_{1}\cdots i_{s}\cdots i_{J}}v_{j_{1}\cdots j_{t}\cdots j_{M}}.$
The TT decomposition serves as a potent alternative to traditional matrix decomposition techniques, which decompose a large tensor into a list of tensor factors [40] by TT-SVD method. As we show in Fig. 1 (a), to decompose the weight matrix into small tensors, it begins by reshaping a matrix $\boldsymbol{W}\in\mathbb{R}^{P\times Q}$ into a higher-dimensional tensor $\mathcal{W}\in\mathbb{R}^{k_{1}\times\cdots\times k_{J}}$. Then $\mathcal{W}$ can be parameterized compactly via a sequence of $J$ tensor factors $\mathcal{G}_{1},\ldots,\mathcal{G}_{J}$ as:
$\mathcal{W}=\mathcal{G}_{1}\times_{3,1}\mathcal{G}_{2}\times_{3,1}\cdots\times_{3,1}\mathcal{G}_{J},$ (1)
Here, each tensor factor $\mathcal{G}_{j}$ has the shape of $\mathcal{G}_{j}\in\mathbb{R}^{r_{j-1}\times k_{j}\times r_{j}}$, where $\boldsymbol{r}=(r_{0},r_{1},\cdots,r_{J})$ is the tensor rank, and the product of dimensions $\Pi_{j}k_{j}=P\cdot Q$. The setup of tensor ranks follows the boundary conditions with $r_{0}=r_{J}=1$, while the other ranks $r_{j},j\notin{0,J}$ are chosen based on the specific tasks or made adaptive [41].
Instead of performing TT decomposition on the weight matrix, we directly initialize, store, and update the list of tensor factors in this work, as shown in Fig. 1 (a). During the forward pass, the tensor factors are directly contracted with the vector of activation values, and the weight matrix $\boldsymbol{W}$ does not need to be reconstructed. Since the size of the tensor factors is small, the contraction process is significantly faster than the original matrix-vector product.
4 Proposed Algorithm
In this section, we introduce the proposed FedTT method, which is built on a novel parameter-efficient tensorized adapter. This adapter is integrated into clients’ local models to effectively adapt LLMs. We begin with discussing the structure and setup of the tensorized adapters, followed by the introduction of FedTT and its enhanced version, FedTT+, which can be selected based on communication budget constraints. The overall workflow of our FedTT method is shown in Fig. 1 (b).
4.1 Tensorized Adapters
To facilitate fine-tuning in our proposed tensorized FL framework, we incorporate tensorized adapters designed to adapt LLMs with minimal additional parameters. These adapters are underpinned by the novel tensorized linear layer [39, 42], which replaces the matrix weight in linear layers with more parameter-efficient TT format weights. This section first details the architecture of the tensorized linear layer and then elucidates the process of constructing tensorized adapters using this innovative layer.
In Sec. 3.2, we introduced how to represent a weight matrix in the TT format, where the TT format tensor factors $\mathcal{G}_{1},\ldots,\mathcal{G}_{J}$ are initialized, stored, and updated during the fine-tuning process. The tensorized layer is successfully integrated into matrix-based LLMs, performing tensor contraction over the list of tensor factors and reshaping them to match the traditional weight matrix $\boldsymbol{W}$. Compared to traditional matrix-based linear layers, the tensorized layers only store the smaller tensor factors, significantly reducing the number of parameters. For instance, a weight matrix of size $16\times 768$, typically containing around $6.1K$ trainable parameters, can be represented by tensor factors totaling approximately $0.6K$ trainable parameters. This dramatic reduction is achieved by assuming a tensor rank of 5 and reshaping $\mathcal{W}$ into dimensions $[4,4,8,8,12]$.
We now introduce the tensorized adapters, built based on the tensorized layers shown in Fig. 1. These adapters utilize a bottleneck structure similar to that described in the original sequential adapter methodology [7], comprising two tensorized linear layers with a nonlinear layer in between. The bottleneck structure further reduces the number of trainable parameters by decreasing the channels connected between the two tensorized layers and the nonlinearity layer, from the size of the hidden dimension to a smaller number, such as 64.
As illustrated in Fig. 1 (b), the tensorized adapter is strategically placed after the attention and MLP components of an encoder/decoder block. In practical applications, we compress the classification head (linear layer) into a tensorized layer and make it trainable for sequence classification tasks, as shown in Fig. 1 (c). Instead, we retain the original language model head without compression for general language modeling tasks, as experiments show a significant reduction in performance when the language modeling head is compressed.
4.2 FedTT Method
Algorithm 1 FedTT
1:for communication round $t\leftarrow 1$ to $T$ do
2:     for clients $c_{i}\in\mathcal{C}$ in parallel do
3:         for local update $k\leftarrow 1$ to $K$ do
4:              $\bm{w}_{i}^{(t)+k+1}\leftarrow\text{Update}(\bm{w}_{i}^{(t)+k},\mathcal{D}_{i})$
5:         end for
6:         Client $c_{i}$ sends $\bm{w}_{i}^{(t)+K}$ to the server
7:     end for
8:     At server: $\bm{w}^{(t+1)}=\frac{1}{N}\sum_{i=1}^{N}\bm{w}_{i}^{(t)+K}$
9:     Server sends $\bm{w}^{(t+1)}$ to all clients in $\mathcal{C}$
10:end for
We introduce the FedTT algorithm, as outlined in Alg. 1. In FedTT, the full model weights and architecture are initially distributed to all clients in the set $\mathcal{C}=\{c_{1},\cdots,c_{N}\}$ at the start of the fine-tuning process. Tensorized adapters are then injected into each client’s local model and designated as the trainable parameters. FedTT operates over $T$ communication rounds, where each client performs $K$ local updates on its trainable parameters during each round.
During communication round $t$, each client $c_{i}\in\mathcal{C}$ updates its trainable parameters $\bm{w}^{(t)}_{i}$, including the tensorized classifier, tensorized adapter, and bias terms, for $K$ local updates. Afterward, each client sends their updated parameters, $\bm{w}^{(t)+K}_{i}$, to the central server. The server then aggregates the weights from all clients as $\bm{w}^{(t+1)}=\sum_{i=1}^{N}\bm{w}^{(t)+K}_{i}$ to obtain the updated global weights for the next training round and sends the aggregated weights $\bm{w}^{(t+1)}$ back to the clients. Note that while we have primarily described cross-silo FL scenarios, the large-scale cross-device FL setting operates similarly, with the key distinction being that in each communication round, the server sends the aggregated weights to a randomly chosen subset of clients $\mathcal{S}\subseteq\mathcal{C}$, rather than to all clients in $\mathcal{C}$.
Unlike most previous Federated PEFT methods such as LoRA, which require transmitting large matrices between clients and the central server, FedTT significantly reduces communication overhead by only transferring small tensor factors. This approach leads to over 10×10$\times$10 × less communication cost compared to LoRA adapters, as demonstrated in Sec. 4.1, making FedTT particularly advantageous in FL scenarios where efficient communication is critical.
4.3 FedTT+ Method
In this section, we propose an improved version of the FedTT method, named FedTT+ to further reduce the number of trainable parameters and enhance FedTT’s suitability for scenarios with data heterogeneity. Before introducing FedTT+, we first explain the intuitive idea behind it.
In FedTT, the loss for back-propagation is computed on the product of tensor factors $\mathcal{G}_{1},\ldots,\mathcal{G}_{J}$. However, under the federated setup, the aggregations in the server are performed separately on $\mathcal{G}_{1},\ldots,\mathcal{G}_{J}$. This practice introduces additional terms in the product of the averaged $\mathcal{G}_{1},\ldots,\mathcal{G}_{J}$, which may slow down the convergence of the algorithms. Ideally, the aggregation on the server should be performed on the product of the low-rank matrices $\mathcal{G}_{1},\ldots,\mathcal{G}_{J}$. The left-hand side of Eq. (2) shows the parameters after aggregation with FedTT using FedAvg, while the right-hand side represents the ideal aggregation.
$\left(\frac{1}{N}\sum_{i=1}^{N}\mathcal{G}_{i,1}\right)\times_{3,1}\ldots\times_{3,1}\left(\frac{1}{N}\sum_{i=1}^{N}\mathcal{G}_{i,J}\right)\neq\frac{1}{N}\sum_{i=1}^{N}\mathcal{G}_{i,1}\times_{3,1}\ldots\times_{3,1}\mathcal{G}_{i,J},$ (2)
where $\mathcal{G}_{i,j}$ is the tensor factor $j$ for clientuser $c_{i}$.
It is important to note that the difference between the right-hand side and the left-hand side of Eq. (2) becomes more significant when: i) the number of clients is large, ii) the clients have non-IID data distributions, and iii) the number of local updates in each communication round is substantial.
In FedTT+, we alleviate this interference problem by freezing most of the tensor factors in each communication round and only updating a small fraction of them. In this approach, most of the tensor factors remain fixed and identical across all clients during training while only a few tensor factors are trainable in each communication round. For example, assume that we just update $\mathcal{G}_{1}$ and freeze $\mathcal{G}_{2},\ldots,\mathcal{G}_{J}$ in communication $t$. Then, $\mathcal{G}_{2},\ldots,\mathcal{G}_{J}$ are identical across the clients and Eq. (2) can be re-written as
$\left(\frac{1}{N}\sum_{i=1}^{N}\mathcal{G}^{(t)}_{i,1}\right)\times_{3,1}\mathcal{G}^{(t-1)}_{2}\times_{3,1}\ldots\times_{3,1}\mathcal{G}_{J}^{(t-1)}=\frac{1}{N}\sum_{i=1}^{N}\left(\mathcal{G}^{(t)}_{i,1}\times_{3,1}\mathcal{G}^{(t-1)}_{2}\times_{3,1}\ldots\times_{3,1}\mathcal{G}_{J}^{(t-1)}\right).$
This modification to FedTT can enhance accuracy in FL scenarios, particularly in cases of severe data heterogeneity. The detailed algorithm is presented in Algorithm 2.
FedTT+ operates similarly to FedTT, but with a key difference: in each communication round $t$, an index $r$ is selected from the set $r\in\{2,\ldots,J-1\}$ (line 3 in Alg. 2). In this process, $\{\mathcal{G}_{1},\mathcal{G}_{r},\mathcal{G}_{J}\}$ are set as trainable parameters (as the first and last tensors are always trained), while the other factors, $\{\mathcal{G}_{2},\ldots,\mathcal{G}_{r-1},\mathcal{G}_{r+1},\ldots,\mathcal{G}_{J-1}\}$, remain frozen (line 5 and 6 in Alg. 2). As a result, clients only send their updated $\{\mathcal{G}_{i,1},\mathcal{G}_{i,r},\mathcal{G}_{i,J}\}$ to the server for aggregation, significantly reducing communication overhead.
[36, 37] have also demonstrated that freezing certain parameters improves the performance of LoRA in the presence of data heterogeneity, particularly with larger models like RoBERTa-large. In the numerical section, we compare FedTT+ with their method under the same settings and show that our approach achieves comparable or better accuracy while using even fewer trainable parameters.
Algorithm 2 FedTT+
1:for communication round $t\leftarrow 1$ to $T$ do
2:     for clients $c_{i}\in\mathcal{C}$ in parallel do
3:         $r\leftarrow$ (choose index $r$ s.t. mod$(t,{\scriptsize J-2})$ $=$ $r-1$)
4:         for local update $k\leftarrow 1$ to $K$ do
5:              for $h\in\{1,r,J\}$:
6:      $\mathcal{G}_{i,h}^{(t)+k+1}\leftarrow\text{Update}(\mathcal{G}_{i,h}^{(t)+k};\mathcal{D}_{i})$
7:         end for
8:         Client $c_{i}$ sends $\{\mathcal{G}_{i,1}^{(t)+K},\mathcal{G}_{i,r}^{(t)+K},\mathcal{G}_{i,J}^{(t)+K}\}$ to server
9:     end for
10:     At server: $\mathcal{G}^{(t+1)}_{h}=\frac{1}{N}\sum_{i=1}^{N}\mathcal{G}_{i,h}^{(t)+K}$ for $h\in\{1,r,J\}$
11:     Server sends $\{\mathcal{G}_{1}^{(t+1)},\mathcal{G}_{r}^{(t+1)},\mathcal{G}_{J}^{(t+1)}\}$ to all clients in $\mathcal{C}$
12:end for
5 Empirical Results
5.1 Experiments Setup
We conducted extensive experiments to evaluate the performance of the proposed algorithms across various language models. For the BERT-family models, we utilized RoBERTa-base [43], DeBERTa-base [44], and RoBERTa-large [43], while for large-scale models, we employed LLaMA-2 [2] (result for LLaMA-2 is presented in Appendix B). Using these models, we compared the proposed algorithm against several PEFT methods in FL scenarios, including BitFit, LoRA, Adapter, Prefix-Tuning, and Prompt-Tuning. Additionally, we benchmarked it against SOTA federated PEFT methods such as FFA-LoRA [36] and RoLoRA [37].
In our experiments, we consider two main FL scenarios: the cross-silo FL scenario [45] and the large-scale cross-device FL scenario [46]. The cross-silo scenario is suitable for networks with typically fewer than $100$ clients. In this case, the server sends the updated model to all clients, meaning every client participates in training during each communication round. In contrast, large-scale cross-device FL is more appropriate for environments with thousands of clients, where only a randomly selected subset of clients is involved in each training round. For cross-silo FL, the number of clients is set to $n\in\{5,10,20,50\}$, while for large-scale cross-device FL, similar to [14], we randomly select $10$ clients from a pool of $1000$. All experiments were conducted using the AdamW optimizer [47], with a similar learning rate and batch size setup across different methods. We performed the experiments on NVIDIA A6000 and V100 GPUs.
Table 1: Comparative analysis of various federated PEFT methods using the DeBERTa-Base model in a cross-silo FL setting with an i.i.d. data distribution and $5$ clients.
Model & Method # Param. MRPC SST-2 QNLI QQP MNLI Avg.
DeBERTa-Base (LoRAr=8) $0.30M$ $91.87$ $94.95$ $92.68$ $89.2$ $87.31$ $91.10$
DeBERTa-Base (P-Tuning) $0.30M$ $82.01$ $90.48$ $82.12$ $84.0$ $80.74$ $83.87$
DeBERTa-Base (LoRAr=4) $0.15M$ $91.72$ $\bm{94.95}$ $\bm{92.66}$ $86.7$ $\bm{86.91}$ $90.58$
DeBERTa-Base (BitFit) $0.10M$ $91.33$ $94.72$ $91.89$ $88.4$ $86.02$ $90.47$
DeBERTa-Base (RoLoRAr=4) $0.08M$ $91.17$ $94.61$ $92.40$ $87.9$ $86.27$ $90.47$
DeBERTa-Base (Prompt) $\bm{0.01M}$ $82.96$ $92.32$ $82.13$ $80.5$ $74.46$ $82.47$
DeBERTa-Base (FedTT) $0.06M$ $\bm{92.68}$ $94.61$ $92.02$ $\bm{88.4}$ $85.99$ $\bm{90.74}$
DeBERTa-Base (FedTT+) $\bm{0.02M}$ $92.60$ $93.58$ $90.54$ $87.9$ $85.33$ $89.99$
Table 2: Comparative analysis of various federated PEFT methods using the RoBERTa-Base model, focusing on cross-silo FL for MRPC and SST-2 tasks and large-scale cross-device FL for other tasks. The reported accuracy for federated LoRA, Adapter, Prefix, and BitFit methods is sourced from [14].
Model & Method # Param. MRPC SST-2 QNLI QQP MNLI Avg.
RoBERTa-Base (Prefix) $3.50M$ $88.1$ $93.7$ $84.6$ $81.8$ $80.4$ $85.7$
RoBERTa-Base (P-Tuning) $0.88M$ $86.8$ $92.1$ $85.6$ $81.5$ $79.8$ $85.2$
RoBERTa-Base (Adapter) $0.70M$ $88.5$ $94.0$ $85.9$ $\bm{87.0}$ $\bm{84.9}$ $88.1$
RoBERTa-Base (IA3) $0.65M$ $88.0$ $93.0$ $\bm{89.4}$ $85.4$ $82.7$ $87.7$
RoBERTa-Base (LoRA) $0.30M$ $\bm{89.8}$ $\bm{94.4}$ $86.0$ $86.5$ $84.7$ $88.3$
RoBERTa-Base (BitFit) $0.10M$ $88.6$ $92.8$ $80.5$ $84.0$ $80.7$ $85.3$
RoBERTa-Base (FedTT) $\bm{0.06M}$ $88.9$ $93.8$ $88.9$ $86.2$ $84.2$ $\bm{88.4}$
5.2 Performance on the BERT Family
We conducted experiments using the Generalized Language Understanding Evaluation (GLUE) benchmark [48], employing the complete training dataset for each task. We recorded the best validation results after the $100$ communication round for cross-silo FL and after the $1000$ communication round for large-scale cross-device FL. Initially, we compared FedTT and FedTT+ with various PEFT methods in a cross-silo FL setting, using the DeBERTa-base model and assuming independent and identically distributed (i.i.d.) data among clients. We set the number of clients to $5$ and the local training epochs to $1$. The results, shown in Table 1, indicate that FedTT achieves a higher average score compared to other methods with lower than $0.15M$ trainable parameters. Notably, FedTT exhibits only a $0.5\%$ performance gap compared to LoRA with a rank of $8$, which has 5×5$\times$5 × more trainable parameters. FedTT+ achieves comparable accuracy to other methods while drastically reducing the number of trainable parameters. It exhibits only a $0.6\%$ performance gap compared to LoRA with a rank of $4$, while using 6×6$\times$6 × fewer trainable parameters, which significantly lowers communication overhead.
We then used the RoBERTa-base model to further evaluate FedTT against other PEFT methods. For the MRPC and SST-2 tasks, we employed a cross-silo FL setup with $10$ clients, while for the other tasks, we applied a large-scale cross-device FL configuration with $1000$ clients, randomly selecting $10$ clients per round. Our settings for the RoBERTa-base model align with those in [14], allowing us to leverage their results of PEFT methods like LoRA, Adapter, Prefix, and BitFit under FL settings. The results in Table 2 show that FedTT achieves a higher average score with the fewest trainable parameters among all methods. Additional details regarding Tables 1 and 2 are provided in Appendices A.3 and A.4, respectively.
5.3 Impact of Data Heterogeneity
Similar to [37], we simulate varying levels of data heterogeneity by using different numbers of clients: $3$ clients (i.i.d. data distribution), $20$ clients (mild heterogeneity), and $50$ clients (severe heterogeneity), with $20$ local updates to further increase data heterogeneity. Additional details are provided in Appendix A.5. Our RoBERTa-Large model settings align with those in [49], allowing us to leverage their results for LoRA, FFA-LoRA, and RoLoRA methods. As shown in Table 3, FedTT+ consistently outperforms these methods across different heterogeneity settings while utilizing fewer parameters. Fig. 2 illustrates test accuracy versus the number of communication rounds in the severe heterogeneity scenario, highlighting the faster convergence of FedTT+ compared to RoLoRA across various tasks.
Table 3: Comparison of SOTA cross-silo FL methods using RoBERTa-Large models under varying degrees of data heterogeneity. The accuracies for LoRA, FFA-LoRA, and RoLoRA methods are sourced from [37].
Data Dist. Model & Method # Param. SST-2 QNLI QQP MNLI Avg.
i.i.d. RoBERTa-Large (LoRAr=2) $68K$ $\bm{95.64}$ $92.04$ $85.85$ $86.16$ $89.92$
RoBERTa-Large (FFA-LoRAr=2) $34K$ $94.91$ $90.11$ $84.06$ $85.48$ $88.64$
RoBERTa-Large (RoLoRAr=2) $34K$ $95.60$ $91.62$ $85.66$ $86.16$ $89.76$
RoBERTa-Large (FedTT+) $\bm{27K}$ $\bm{95.64}$ $\bm{94.05}$ $\bm{87.99}$ $\bm{87.27}$ $\bm{91.24}$
mild het. RoBERTa-Large (LoRAr=2) $68K$ $94.27$ $86.91$ $81.22$ $82.07$ $86.12$
RoBERTa-Large (FFA-LoRAr=2) $34K$ $93.92$ $89.58$ $80.51$ $82.62$ $86.66$
RoBERTa-Large (RoLoRAr=2) $34K$ $94.84$ $90.77$ $85.13$ $85.10$ $88.96$
RoBERTa-Large (FedTT+) $\bm{27K}$ $\bm{95.64}$ $\bm{92.60}$ $\bm{87.46}$ $\bm{87.69}$ $\bm{90.85}$
sever het. RoBERTa-Large (LoRAr=2) $68K$ $93.23$ $82.57$ $58.96$ $76.96$ $77.93$
RoBERTa-Large (FFA-LoRAr=2) $34K$ $92.32$ $85.15$ $62.79$ $77.78$ $79.51$
RoBERTa-Large (RoLoRAr=2) $34K$ $\bm{94.61}$ $89.83$ $85.15$ $85.55$ $88.78$
RoBERTa-Large (FedTT+) $\bm{27K}$ $94.50$ $\bm{90.17}$ $\bm{86.65}$ $\bm{86.28}$ $\bm{89.40}$
Figure 2: Test accuracies versus communication rounds for RoBERTa-Large models across different tasks in a cross-silo FL setting with 50 clients.
6 Conclusion
This work presents a novel federated and tensorized fine-tuning framework, addressing the key challenges of communication overhead and data heterogeneity in FL. FedTT leverages tensorized adapters to significantly reduce communication costs while maintaining high performance, achieving up to a 10×10$\times$10 × reduction in trainable parameters compared to existing methods. Additionally, FedTT+ enhances robustness in cross-silo FL by adaptively freezing portions of tensor factors, further optimizing parameter efficiency. Extensive experiments with models like BERT and LLaMA-2 demonstrate the effectiveness of FedTT and FedTT+ in both cross-silo and large-scale cross-device FL settings, offering a scalable and communication-efficient solution for fine-tuning large language models in distributed environments.
Limitations
In this work, while we addressed the data heterogeneity challenge, we did not conduct experiments on another significant challenge in FL: system heterogeneity. However, our framework has the potential to tackle this challenge as well, by allowing different tensor ranks to be assigned to clients based on their computational capabilities. This presents an interesting direction for future research.
Although the primary motivation for using FL is to enhance data privacy, different algorithms provide varying levels of privacy protection, as privacy leakage is inevitable in most methods. A key experiment missing from our work is a comparison of our algorithm with others in terms of privacy preservation. Given that a lower number of trainable parameters could intuitively lead to better privacy in FL settings, our framework is likely to offer strong privacy-preserving capabilities, which presents an interesting area for future investigation.
Appendix A Experiment setup
A.1 Dataset Setup
We began our experiments using the Generalized Language Understanding Evaluation (GLUE) benchmark [48], which includes a range of natural language understanding tasks. These tasks consist of perceptual analysis (SST2 [50]), similarity and paraphrase tasks (MRPC, QQP [51]), and natural language reasoning (MNLI, QNLI [52, 53]). The utilized metrics for the GLUE benchmark are summarized in Table 4. We recorded the best validation results after $100$ communication rounds for cross-silo FL and after $1000$ communication rounds for large-scale cross-device FL. For the DeBERTa-Base models on the QQP task, we randomly selected $1000$ samples from the validation set and reported the highest accuracy achieved.
Table 4: Dataset descriptions and statistics.
Task # Train # Dev. Metric
MRPC 3,301 367 F1 Score
SST-2 66,675 674 Accuracy
QNLI 103,695 5,463 Accuracy
QQP 360,210 40,430 Accuracy
MNLI 388,774 9,815 Accuracy
We then selected two multiple-choice tasks (COPA and ReCoRD) from the SuperGLUE benchmark [54] and a question-answering generation task (SQuAD [55]). The metrics used for evaluation are summarized in Table 5.
Table 5: The utilized metrics for the SuperGLUE benchmark.
Task Name Metric
COPA Accuracy
ReCoRD F1
SQuAD F1
A.2 Additional Detail of TT-format
In this paper, we use the TT format to structure the weight matrices within the tensorized layers. To accommodate matrices of varying shapes, we design specific tensor shapes based on the hidden sizes and bottleneck configurations of different models. The tensor shapes are outlined in Table 6, with examples provided for the DeBERTa/RoBERTa-base and LLaMA-2-7b models, which have hidden sizes of 768 and 4096, respectively. For models with other hidden sizes, the appropriate tensor shape must be determined prior to training.
Table 6: The shape settings of the TT-format
Modules Matrix Shape Tensor Shape
Tensorized Adapters $768\times 64$ [8, 8, 12, 8, 8]
$4096\times 64$ [16, 16, 16, 4, 4, 4]
$64\times 768$ [8, 8, 12, 8, 8]
$64\times 4096$ [4, 4, 4, 16, 16, 16]
Tenosrized Classifier (Optional) $768\times 768$ [12, 8, 8, 8, 8, 12]
$768\times 768$ [8, 8, 8, 8, 8, 8, 8, 8]
A.3 Additional Details for Table 1
We use DeBERTa-Base models with the GLUE dataset in a cross-silo FL setup. The learning rate is selected from $[5e-3,1e-3,5e-4,1e-4]$, and the batch size from $[16,32]$ for all tasks and methods. The number of clients is set to $5$, with one local update per communication round. The best validation results were recorded after $100$ communication rounds. Additional details on the chosen parameters are provided in Table 7.
Table 7: The hyperparameter grids used for GLUE experiments.
Experiment Hyperparameters Values
LoRA Rank ${4,8}$
RoLoRA Rank $4$
Bitfit Bias Terms All
Prompt # Tokens 10
P-tuning # Tokens 20
Prompt Length $[128,768]$
FedTT Tensor Rank $5$
FedTT+ Tensor Rank $5$
Table 8: Number of trainable parameters for results in Table 3.
Model & Method SST-2 QNLI QQP MNLI Avg.
RoBERTa-Large (LoRAr=2) $49K$ $74K$ $74K$ $74K$ $68K$
RoBERTa-Large (FFA-LoRAr=2) $\bm{25K}$ $37K$ $37K$ $37K$ $34K$
RoBERTa-Large (RoLoRAr=2) $\bm{25K}$ $37K$ $37K$ $37K$ $34K$
RoBERTa-Large (FedTT+) $27K$ $\bm{27K}$ $\bm{27K}$ $\bm{27K}$ $\bm{27K}$
Table 9: Comparative analysis of LoRA, FedTT, and FedTT+ using the LLaMA2-7B and LLaMA2-13B models.
Fed. Set. Model & Method # Param. COPA ReCoRD SQuAD Avg.
large-scale FL LLaMA2-7B (LoRAr=8) $4.19M$ $87$ $81.0$ $90.49$ $86.16$
LLaMA2-7B (FedTT) $0.52M$ $90$ $80.1$ $89.32$ $86.47$
cross-silo FL LLaMA2-13B (LoRAr=8) $6.55M$ $89$ $83.6$ $91.07$ $87.89$
LLaMA2-13B (FedTT) $0.64M$ $89$ $83.5$ $90.06$ $87.52$
LLaMA2-13B (FedTT+) $0.18M$ $88$ $83.1$ $89.12$ $86.74$
A.4 Additional Details for Table 2
We follow the experimental setup of [14], using RoBERTa-Base models with the GLUE dataset in both cross-silo and large-scale cross-device FL configurations. The learning rate was selected from $[1e-2,5e-3,1e-3,5e-4,1e-4,5e-5]$, and batch size from $[16,32]$ across all tasks and methods. For the MRPC and SST-2 tasks, we used a cross-silo FL setup with $10$ clients, while for the other tasks, we employed a large-scale cross-device FL setup with $1000$ clients, selecting $10$ clients per round. Accuracy results for Prefix, Adapter, LoRA, and BitFit methods are sourced from [14]. In addition to reporting the accuracy of the FedTT method, we also included the P-Tuning method with prompt lengths $[128,768]$, and IA3 [56] method, to further enrich the experiments.
A.5 Additional Details for Table 3
We adopt the same settings as [37], using RoBERTa-Large models with the GLUE dataset in cross-silo FL scenarios. To simulate varying levels of data heterogeneity, we report the accuracy of our method for $3$ clients (i.i.d. data distribution), $20$ clients (moderate heterogeneity), and $50$ clients (high heterogeneity), with $20$ local updates. The learning rate is set to $31e-$ 3 for all tasks, with a batch size of $64$ for SST-2 and $32$ for other tasks. In [37], to reduce the number of trainable parameters, LoRA is applied selectively to certain layers: the last $6$ layers for SST-2, and the last $9$ layers for QNLI, QQP, and MNLI. Consequently, the number of trainable parameters varies depending on the task. We provide a comparison of the number of trainable parameters for LoRA, FFA-LoRA, RoLoRA, and our proposed FedTT+ method in Table 8. Accuracy results for LoRA, FFA-LoRA, and RoLoRA are sourced from [37].
Appendix B Performance Comparison on the Large-Scale Language Models
Due to the large number of parameters in LLaMA-2, it is rarely used in FL scenarios. None of the existing works on LLMs provide results for such models. However, by employing FedTT, the significantly reduced number of communication overhead allows us to utilize these models. We use SuperGLUE tasks [54] and a generation task (SQuAD [55]) to compare FedTT and FedTT+ with LoRA. The results are shown in Table 9. We used the LLaMA2-7B model to simulate large-scale cross-silo FL with $1000$ clients, where $10$ of them are chosen randomly in each round for training. FedTT has performance nearly identical to LoRA with a rank of $8$, while achieving about 10×10$\times$10 × lower communication overhead. For cross-silo FL, we utilized the LLaMA2-13B model with $10$ clients. As seen in Table 9, both FedTT and FedTT+ demonstrate performance close to LoRA, with 10×10$\times$10 × and 30×30$\times$30 × lower communication overhead, respectively.